# Semi-orthogonality in Fukaya-Seidel mirrors to blowups of abelian varieties

### v1

@@ 3.2.4. Inverses: continuation maps
$\frac{\partial(\theta\circ\tilde{u})}{\partial s^{UHP}}\mid_{(s_{0},t_{0})}>0.$
In words, the image $\tilde{u}$ moves counterclockwise in the base $\mathbb{C}^{*}$ as we move in the direction of increasing $s$ on the boundary of the upper-half plane in the domain. From the definition of $\tilde{u}$, this movement has boundary on a fixed Lagrangian, either $W_{0}(L_{0}^{k})$ or $W_{0}(L_{1}^{j})$ depending on which connected component $(s_{0},t_{0})$ is on. This Lagrangian has ends which have strictly increasing radial coordinate outside the disc $\{|w|\leq r_{2}\}$ by Definitions 3.4 and 3.6. The tangent to the curve on the Lagrangian then has a positive radial component in the $s$ direction. Now we compare the two directions. Our analysis above showed that the curve $\tilde{u}$ has a positive angular component and zero radial component in the $s$ direction at $(s_{0},t_{0})$ which means it is rotated 90 degrees counterclockwise from the radial direction, and hence also counterclockwise from the Lagrangian boundary tangent vector at the same point. But, the isotopy from $\tau$ moves the Lagrangian clockwise. Since the two are going in opposite directions, this is a contradiction, and the curve $\tilde{u}$ could not achieve its maximum on the boundary outside $\tilde{A}$. For boundedness away from 0, we apply the above argument to $1/W_{0}$ on the inverted annulus $1/\tilde{A}$ and use Equation (3.10). ∎
2b) Now we exclude the possible types of bubbling described in Gromov compactness. We have the assumptions needed for Gromov compactness to hold by [AA24, Proposition 3.13] where their $X_{H}$ and $H$ terms are 0 in our setting since we don’t have wrapping in the fibers. The assumptions of their lemma hold because we proved in Lemma 3.13 above that $[u]$ is contained in a compact subset which we labeled $\Omega$ to match their notation.
Possible degenerations for discs are sphere bubbles (energy concentrates at an interior point), disc bubbles (energy concentrates at a boundary point), or strip-breaking (energy concentrates at a limiting intersection point). We can exclude sphere bubbles similar to [Can20, Corollary 4.53], because discs are somewhere injective by the paragraph above Lemma 3.12. Since Lagrangians $L$ are homeomorphic to $T^{2}$ and $\pi_{2}(T^{2})=0$, a disc bubble $u$ would have $u]\in\pi_{2}(T^{2},\rho^{t}(L))=0$ so must be constant. Therefore there can be no disc bubbles. Strip-breaking, from reparametrization by translation, is excluded because the prescribed moving Lagrangian boundary condition prevents translation as an automorphism; the continuation maps count index 0 strips which limit to fixed Lagrangian boundaries. So the moduli space cannot limit to any configurations not already in it and therefore is compact.
This concludes the proof of Theorem 3.10 that the moduli space $\mathcal{M}(p,q;[u],J)$ can be counted.
@@
Taking them to represent elements in the quotients $\text{colim}_{j\to\infty}HF^{*}(L_{0}^{k},L_{1}^{j})=\coprod_{j}HF^{*}(L_{0}^{k},L_{1}^{j})/\sim$ and $\text{colim}_{j\to\infty}HF^{*}(L_{0}^{k+1},L_{1}^{j})=\coprod_{j}HF^{*}(L_{0}^{k+1},L_{1}^{j})/\sim$ respectively, the induced continuation maps on cohomology $F_{L_{0}^{k},L_{1}^{j}}:HF(L_{0}^{k},L_{1}^{j})\to HF(L_{0}^{k+1},L_{1}^{j+1})$ form an inverse to
$\mu^{2}(-,e_{L_{0}}^{k}):\text{colim}_{j\to\infty}HF^{*}(L_{0}^{k+1},L_{1}^{j})\to\text{colim}_{j\to\infty}HF^{*}(L_{0}^{k},L_{1}^{j}).$
The proof uses [AA24, Lemma 3.19] that Equation (3.27) holds up to homotopy on the chain-level by constructing parametrized moduli spaces [Sei08, p. 241], [Can20, Lemma 5.1] varying choices in 1 dimensional families and taking the boundary. It still applies in our setting because we have a maximum principle.
@@ 3.2.5. Morphisms in the derived category
We would like to glue together equivalent objects. A motivating example [Wei94, §1.5] is that of making a simplicial complex $X$ contractible by gluing it to a point. Taking its cone $CX$ then $C_{j}(CK)=C_{j-1}(K)\oplus C_{j}(K),\;j\geq 1$ from chains fully in the base versus those through the cone vertex $s$. Checking the differential and quotienting by $C_{0}$ of the cone point, this is just $C_{j}(CK)/C_{j}(\text{pt})=Cone_{j}(\mathbf{1}_{*}:C_{\bullet}(K)\to C_{\bullet}(K))$. So the notion of topological and algebraic cone coincide for topological spaces.
In practice when computing morphisms of Lagrangians in a Landau-Ginzburg model, the informal way of doing so is to “push the ends of the input Lagrangian, when projected to the base, to lie above the ends of the output Lagrangian.” In other words, if $L_{0}^{+}$ is $L_{0}$ with the ends pushed up, we would like a morphism $L_{0}\to L_{1}$ to be a “roof” $L_{0}^{+}$$L_{0}$$L_{1}$$\scriptstyle{e}$$\scriptstyle{p}$ where $e\in CF(L_{0}^{+},L_{0})$ is a quasi-unit and $p\in L_{0}^{+}\cap L_{1}$ is an intersection point or morphism of $CF(L_{0}^{+},L_{1})$. This roof is the standard construction of a left fraction in [Wei94, Prototype 10.3.5], which is a morphism in a localized category. To compose roofs would involve a pullback along $L_{0}^{+}\rightarrow L_{1}\leftarrow L_{1}^{+}$ [Wei94, Definition 10.3.4, Ore condition]. However, it is not clear what this pullback would be in practice.
We instead take a homotopy colimit over Yoneda modules $\mathcal{Y}_{L^{k}}=\mathcal{O}(-,L^{k})$ so we are in the abelian category of $\mathcal{O}$-modules. We follow [AA24, §3.5], where they use standard constructions similar to [AS10, §(3g)], [GPS20, §3.1.3]. These refer to [LO06] for localization of $A_{\infty}$-categories, a process that generalizes localization of categories [Wei94, §10.3], [KS06, §10.2], [Nee01, §1], [Rie14, Example 11.5.11], [BN93].
@@
$\mathcal{Y}_{L^{\infty}}:=\text{hocolim}_{\mathcal{O}}\mathcal{Y}_{L^{k}}$
over the directed category $\mathcal{O}$ and morphisms
@@
So instead of Floer theory with $L$, we use $\mathcal{Y}_{L^{\infty}}$. This makes sense as the right object to use because we have a sequential diagram
$\ldots\mathcal{Y}_{L^{-1}}\to\mathcal{Y}_{L^{0}}\to\mathcal{Y}_{L^{1}}\to\ldots$
with connecting maps given by multiplication of quasi-units $\mu^{2}(e_{L^{k}},-):\mathcal{Y}_{L^{k}}\to\mathcal{Y}_{L^{k+1}}$. The homotopy colimit identifies objects that have a homotopy between them, preserving weak equivalences which are when a map of spaces induces an isomorphism on homotopy groups (so a particular inverse isn’t needed, as for a strong equivalence). This is similar to how a colimit glues along connecting maps, but a limit such as the pullback of $0\hookrightarrow[0,1]\hookleftarrow 1$ is empty even though $0,1,[0,1]$ are all homotopy equivalent, while the homotopy limit would be $\{\gamma:[0,1]\to[0,1]\mid\gamma(0)=0,\gamma(1)=1\}$, [Rie14, Exercise 6.5.4]. In the case of Lagrangians in the Fukaya category, that homotopy can be realized geometrically as an isotopy from one Lagrangian to another that more fully wraps. Furthermore, in HMS we take the split-closed derived (also called idempotent completion) of the Fukaya category, [Nee01, Proposition 1.6.8].
Now we justify why this definition for morphisms means we can push the ends of the input Lagrangian up. The following isomorphism implies that after taking cohomology of the chain level construction of $\mathcal{W}(L_{0},L_{1})$, we can take a sufficiently large wrapping of $k$ on the input Lagrangian in usual Floer theory.
@@
(3.29) $\text{colim}_{j\to\infty}HF^{*}(L_{0}^{k},L_{1}^{j})=\coprod_{j\in\mathbb{N}}HF^{*}(L_{0}^{k},L_{1}^{j})/(p\sim\mu^{2}(e_{L_{1}^{j}},p))\to H\mathcal{W}(L_{0},L_{1}).$
This holds because the following is a model for the homotopy colimit.
@@
This is a standard model from topology called the mapping telescope, adapted to abelian categories as described in [AS10, §(3g)], [BN93]. One can use it as the definition as in [Nee01, Definition 1.6.4]. More recent homotopical algebra developments, for example [Rie14, Example 11.5.11], explain why this agrees with the general definition of homotopy colimit as the left derived functor of the colimit constructed as the geometric realization of the cobar construction [Rie14, Corollary 5.1.3]. (A motivating example is the double mapping cylinder [Rie14, Example 6.4.5].)
@@
The proof follows again from homological algebra, the definition of $\mathcal{O}$, and Equation (3.27). These all still apply in our setting. Since the bonding maps in the holim of Result 1 are isomorphisms on cohomology by Result 2, taking a projection we find that for any $k$, Equation (3.29) holds.
3.2.6. Calculations
Here we take $j>i+1$. The results can be extended to all $i,j$ by Serre duality on the A- and B-side [ACLL23a, Lemma 4.7].
Lemma 3.20 (Morphisms on $\mathcal{F}_{1}$).

### v2

@@ 3.2.4. Inverses: continuation maps
(3.37) $\frac{\partial(\theta\circ\tilde{u})}{\partial s^{UHP}}\mid_{(s_{0},t_{0})}>0.$
[…]
Proof of Lemma 3.11.
We exclude the possible types of bubbling described in Gromov compactness in Proposition 3.14. We have the assumptions needed for Gromov compactness to hold by [AA24, Proposition 3.13] where their $X_{H}$ and $H$ terms are 0 in our setting since we don’t have wrapping in the fibers. The assumptions of their lemma hold because we proved in Lemma 3.15 above that $[u]$ is contained in a compact subset which we labeled $\Omega$ to match their notation.
We can exclude sphere bubbles by the dimension argument [Ca20, Corollary 4.53], because discs are somewhere injective by Lemma 3.13. As $W_{0}(L)$ is a U-shape or arc, it cannot bound a disc in the base. And in a fiber, a linear Lagrangian cannot bound discs so $u]\in\pi_{2}(T^{4},\ell_{k})=0$ must be constant. Therefore there can be no disc bubbles. Strip-breaking, from reparametrization by translation, is excluded because the prescribed moving Lagrangian boundary condition prevents translation as an automorphism; the continuation maps count index 0 strips which limit to fixed Lagrangian boundaries. So the moduli space cannot limit to any configurations not already in it and therefore is compact.
This concludes the proof of Lemma 3.11 that the moduli space $\mathcal{M}(p,q;[u],J)$ is compact.
@@
More specifically, taking $p_{k}$ and $p_{k+1}$ to represent elements in the quotients $\underset{j\to\infty}{\text{colim }}HF^{*}(L_{0}^{k},L_{1}^{j})=$
$\coprod_{j}HF^{*}(L_{0}^{k},L_{1}^{j})/\sim$ and $\underset{j\to\infty}{\text{colim }}HF^{*}(L_{0}^{k+1},L_{1}^{j})=\coprod_{j}HF^{*}(L_{0}^{k+1},L_{1}^{j})/\sim$ respectively, the induced continuation maps on cohomology $F_{L_{0}^{k},L_{1}^{j}}:HF(L_{0}^{k},L_{1}^{j})\to HF(L_{0}^{k+1},L_{1}^{j+1})$ form an inverse to
$\mu^{2}(-,e_{L_{0}}^{k}):\underset{j\to\infty}{\text{colim }}HF^{*}(L_{0}^{k+1},L_{1}^{j})\to\underset{j\to\infty}{\text{colim }}HF^{*}(L_{0}^{k},L_{1}^{j}).$
Remark 3.17.
The proof uses [AA24, Lemma 3.19] that Equation (3.38) holds up to homotopy on the chain-level by constructing parametrized moduli spaces [seidel, p. 241], [Ca20, Lemma 5.1] varying choices in 1 dimensional families and taking the boundary. It still applies in our setting because we have a maximum principle.
@@ 3.2.5. Morphisms in the derived category
As motivated at the start of this chapter we take the following
@@
(3.39) $\mathcal{Y}_{L^{\infty}}:=\underset{{k\to\infty}}{\text{hocolim}}\;\mathcal{Y}_{L^{k}}$
over the sequential diagram
(3.40) $\ldots\mathcal{Y}_{L^{k-1}}\xrightarrow[]{\mu^{2}(e_{L^{k-1}},-)}\mathcal{Y}_{L^{k}}\xrightarrow[]{\mu^{2}(e_{L^{k}},-)}\mathcal{Y}_{L^{k+1}}\to\ldots.$
Define morphisms
@@
So instead of Floer theory with $L$, we use $\mathcal{Y}_{L^{\infty}}$. Now we justify why this definition for morphisms means we can push the ends of the input Lagrangian up. The following isomorphism implies that after taking cohomology of the chain level construction of $\mathcal{W}(L_{0},L_{1})$, we can take a sufficiently large wrapping of $k$ on the input Lagrangian, then take the usual Floer theory of Equation (3.18).
@@
(3.42) $\underset{j\to\infty}{\text{colim}}HF^{*}(L_{0}^{k},L_{1}^{j})=\coprod_{j\in\mathbb{N}}HF^{*}(L_{0}^{k},L_{1}^{j})/(p\sim\mu^{2}(e_{L_{1}^{j}},p))\to H\mathcal{W}(L_{0},L_{1}).$
The first step to justify this is true is to note that the following is a model for the homotopy colimit.
@@
This is a standard model from topology called the mapping telescope, adapted to abelian categories as described in [AS10, §(3g)], [BN93]. One can use it as the definition of the homotopy colimit of a sequential diagram as in [Ne01, Definition 1.6.4]. More recent homotopical algebra developments, for example [Ri14, Example 11.5.11], explain why this agrees with the general definition of homotopy colimit as the left derived functor of the colimit constructed as the geometric realization of the cobar construction [Ri14, Corollary 5.1.3]. (A motivating example is the double mapping cylinder [Ri14, Example 6.4.5].)
@@
The proof follows again from homological algebra, the definition of $\mathcal{O}$, and Equation (3.38). These all still apply in our setting. Since the bonding maps in the holim of Claim 3.21 are isomorphisms on cohomology by Lemma 3.22, taking a projection we find that for any $k$, Equation (3.42) holds.
3.3. Lagrangian gradings
Before we compute the morphisms we need to discuss Lagrangian gradings. Here we assume $j>i+1$ so $CF^{*}(\ell_{i},\ell_{j})$ and $CF^{*}(\ell_{i+1},\ell_{j})$ are supported in degree 0 by [ACLL_Bfield, Example 4.6], where $\ell_{k}$ is linear of slope $k$ and the Lagrangian grading is $\frac{1}{\pi}\arctan(k)$. The calculations for the morphisms in Subsection 3.4 can be extended to other $i,j$ using Serre duality on the A- and B-side [ACLL_Bfield, Lemma 4.7], done in the proof of Corollary 4.3.
We start with the standard grading in $\mathbb{C}$. The typical gradings in a bigon between curves $l_{1},l_{2}$ in $\mathbb{R}^{2}$ have degree 0 on the right intersection point and degree 1 on the left intersection point, as in Figure 4 and [fuk_intro, Example 1.9]. On a surface, a grading on a Lagrangian $l_{j}$ is a continuous choice of angle $\alpha_{j}(t)$ along the Lagrangian for the tangent vector field $\dot{l}_{j}(t)$, relative a line field. Then intersection points are graded as follows.
[…]
Figure 4. Angles depicted as fractions of $2\pi$ relative the positive horizontal line field
  • Step 1:
    Find $\alpha_{2}(t)-\alpha_{1}(t)$ at values of $t$ where the paths intersect (direction of the arrows on the base Lagrangians matter).
  • Step 2:
    At that same value of $t$, we rotate the line spanned by $\dot{l}_{2}(t)$ counterclockwise until it coincides with that spanned by $\dot{l}_{1}(t)$ (direction of the arrows on the base Lagrangians don’t matter). Since counterclockwise is always positive, we add that amount of rotation to $\alpha_{2}(t)-\alpha_{1}(t)$. This is referred to as the reverse canonical short path.
  • Step 3:
    Twice this winding number is the degree, because rotating a line back to itself, through an angle $\pi$, should count as one rotation.
In an LG model, fiber and base contributions add, both for Lagrangian gradings and degrees of intersection points, by [ACLL_LG, Lemma 3.11 and Corollary 3.12]. To grade Lagrangians, a choice of holomorphic volume form $\Omega$ on $Y^{0}$ must be made, so locally
(3.44) $\Omega=a(z_{1},z_{2},z_{3})dz_{1}\wedge dz_{2}\wedge dz_{3}$
for local holomorphic coordinates $z_{1},z_{2},z_{3}$ and non-vanishing $a(z_{1},z_{2},z_{3})$. From there, we define a quadratic complex volume form $\Theta:=(\Omega/|\Omega|)^{\otimes 2}$ and this then defines a squared phase map $\alpha_{\Theta}:LGr(TY^{0},\omega)\to U(1)$ by
(3.45) $\alpha_{\Theta}(p,V)=\frac{\Omega(e_{1},e_{2},e_{3})^{2}}{|\Omega(e_{1},e_{2},e_{3})|^{2}}$
where $p\in Y^{0}$, $V=T_{p}L\cong\mathbb{R}^{3}$ is a linear Lagrangian subspace of $(T_{p}Y^{0},\omega(p))$, and $(e_{1},e_{2},e_{3})$ is an ordered $\mathbb{R}$-basis of $V$.
Taking $e_{1},e_{2}$ to span the tangent space of the fiber Lagrangian $\ell_{k}$, we find its squared phase function is $e^{-2\pi i\cdot 2\phi_{k}}$ where $\phi_{k}=\frac{1}{\pi}\arctan(k)$. Its squared phase map is the square of the determinant in coordinates $\left(\begin{matrix}r_{1}\\ r_{2}\end{matrix}\right)=\left(\begin{matrix}2&1\\ 1&2\end{matrix}\right)^{-1}\left(\begin{matrix}\xi_{1}\\ \xi_{2}\end{matrix}\right)$, [ACLL_Bfield, Example 4.6]. On the base take $e_{3}=\frac{\dot{\gamma}_{L}(t)}{|\dot{\gamma}_{L}(t)|}$ to be the unit tangent to the projected curve in the base. The choice of $\Omega$ in the base is then equivalent to a choice of line field with respect to which we measure angles of the tangent vector field $\dot{\gamma}_{L}$ from, as mentioned above for surfaces. We interpolate between two models, a positive horizontal line field away from 0 in $\mathbb{C}^{*}$, and an angular line field around 0 in $\mathbb{C}^{*}$ (positive horizontal line field in the cylinder).
Definition 3.23 (Definition of Lagrangian grading).
Fiber: the relative squared phase map is
$\alpha_{\Theta^{\text{Vert}}}:LGr=T^{4}\times U(2)/O(2)\mapsto U(1),\qquad(x,aO(2))\mapsto\det(a)^{2}.$
Base: For $|w|>R_{2}$, $L=L_{j}$: Note that $(\exp(2\pi i\alpha_{j}(t)))^{2}=\alpha_{L}^{\text{Hor}}(p)$ the horizontal squared phase function of $L_{j}$ of [ACLL_LG, Lemma 3.11, Eq. (3.8)]. For horizontal lift $(\dot{\gamma}_{L_{j}}(t))^{\#}$ and recalling $r(t)$ and $\theta(t)$ parametrize $L_{j}$ as in Definition 3.3, define
For the part of the U-shape around the critical value we take a circle $\dot{r}(t)=0$, $\dot{\theta}(t)=1$, then
so we can take the choice for $\alpha_{j}(t)$ to be
(3.47) $\alpha_{j}(t)=\theta(t)+1/4.$
In other words, $\alpha_{j}(t)$ is the angle that the tangent vector $\dot{\gamma}_{L_{j}}(t)$ makes with respect to the horizontal line field. So we recover the fact that the tangent to the circle is rotated $90$ degrees from the radial vector, when measured with respect to the positive horizontal vector field. In particular, the phase function is 1/4-1/4 at $b_{S}$ by Equation (3.2). For the ends of the U-shape we have $r(t)=R_{4}^{\prime}|t|$ for constant of proportionality $R_{4}^{\prime}$ in Equation (3.2), $\dot{\theta}(t)=0$, and $\theta(t)=\theta_{h}(L)$ so
$e^{2\pi\mathbf{i}\cdot 2\theta_{h}(L)}=e^{2\pi\mathbf{i}\cdot 2\alpha_{j}(t)}$
and we can take the choice for $\alpha_{j}(t)$ to be the choice of angle in $(-\frac{1}{2},\frac{1}{2})$
(3.48) $\alpha_{j}(t)=\begin{cases}\theta_{h}(L)-\lfloor\theta_{h}(L)\rfloor&\text{ if }<\frac{1}{2}\\ \theta_{h}(L)-\lfloor\theta_{h}(L)\rfloor-1&\text{ if }>\frac{1}{2}\end{cases}.$
This is again the angle the tangent vector $\dot{\gamma}_{L_{j}}(t)$ makes with respect to the horizontal line field.
For $|w|>R_{2}$, $L=K_{j}$: For mirror symmetry reasons to be explained below, we take negative the above choice of horizontal squared phase function
(3.49) $\alpha_{K}^{\text{Hor}}(p)=-\alpha_{L}^{\text{Hor}}(p)$
so the phase function is shifted by 1/2-1/2
(3.50) $\alpha_{j}(t)=\begin{cases}\theta_{h}(L)-\lfloor\theta_{h}(L)\rfloor-\frac{1}{2}&\text{ if }<\frac{1}{2}\\ \theta_{h}(L)-\lfloor\theta_{h}(L)\rfloor-\frac{3}{2}&\text{ if }>\frac{1}{2}\end{cases}.$
For $|w|<R_{1}$, $L=K_{j}$: Note that only $W_{0}(K_{j})$ passes through this region, and not $W_{0}(L_{j})$. Here we take
(3.51) $\alpha_{K}^{\text{Hor}}(p)=\frac{(d\log W_{0}(\dot{\gamma}_{K_{j}}(t)^{\#}))^{2}}{|d\log W_{0}(\dot{\gamma}_{K_{j}}(t)^{\#})|^{2}}=\frac{(\dot{\gamma}_{K_{j}}(t))^{2}}{|\dot{\gamma}_{K_{j}}(t)|^{2}}\bigg/\frac{\gamma_{K_{j}}(t)^{2}}{|\gamma_{K_{j}}(t)|^{2}}.$
Without wrapping, $W_{0}(K_{j})$ can be parametrized by $\gamma_{K_{j}}(t)=te^{2\pi\mathbf{i}\theta_{h}(K_{j})}$ and the standard convention would be $\alpha_{j}(t)=0$. We instead take $\alpha_{j}(t)=-1$ to be consistent with $|w|>R_{2}$. With wrapping, by Remark 3.7, instead of a colimit we can consider a quadratically-wrapped Lagrangian around 0 with $H(\mu,\theta)=\frac{1}{2}\mu^{2}$. Then under the time-1 flow of the corresponding Hamiltonian vector field $\mu\frac{\partial}{\partial\theta}$, we find that near 0, $W_{0}(K_{j})$ wraps. The new curve is $\{(\mu,\theta(\mu))\mid\theta(\mu)=\theta_{h}(K_{j})+\mu,\;\mu\in\mathbb{R}_{+}\}$ in action-angle coordinates. In coordinate $w$ on $\mathbb{C}^{*}$, this can be parametrized as $\gamma(t)=te^{\mathbf{i}(2\pi\theta_{h}(K_{j})+\log t)}$, by taking $t=r$ and using that the action coordinate $\mu=\frac{1}{2\pi}\log r$. Thus
(3.52) $\alpha_{K}^{\text{Hor}}(p)$ $=\frac{(\dot{\gamma}_{K_{j}}(t))^{2}}{|\dot{\gamma}_{K_{j}}(t)|^{2}}\bigg/\frac{\gamma_{K_{j}}(t)^{2}}{|\gamma_{K_{j}}(t)|^{2}}=\frac{(1+\mathbf{i})^{2}e^{\mathbf{i}\cdot 2(2\pi\theta_{h}(K_{j})+\log t)}}{|1+\mathbf{i}|^{2}}\bigg/e^{\mathbf{i}\cdot 2(2\pi\theta_{h}(K_{j})+\log t)}$
$=\mathbf{i}=e^{2\pi\mathbf{i}\cdot 2\alpha_{j}(t)}.$
So one choice is $\alpha_{j}(t)=1/8$ but we again shift by 1-1 to be consistent so take $\alpha_{j}(t)=1/8-1=-7/8$. In words, this measures that the universal cover of a curve wrapping around the 0 end of the cylinder is a curve of slope 1 that makes an angle $\pi/4$ with the horizontal line field in $(\mu,\theta)$ coordinates. In the complex plane in the base, it wraps around 0.
For $R_{1}<|w|<R_{2}$, $L=K_{j}$: Again only $W_{0}(K_{j})$ passes through this region. We interpolate between the two choices for $\alpha_{K}^{\text{Hor}}$. Visually, interpolating between the two different choices of horizontal holomorphic one form, $dw$ and $d\log w$, is as follows. Draw a line field on the base, with respect to which tangent vectors of the projected Lagrangians are measured. Outside of the disc of radius $R_{2}$ it is the horizontal line field and inside the disc of radius $R_{1}$ it spirals around the origin. In between, rotate the horizontal lines to the angular ones.
Summary of the base gradings: The Lagrangian U-shapes in the base $W_{0}(L_{j})$ are all graded the same, with phase 1/4-1/4 at $b_{S}$. The grading on $W_{0}(K_{j})$ is shifted by 1-1 from that of the ends of the U-shapes because the phase is shifted by 1/2-1/2. This is done to match morphisms on the B-side in the non-orthogonal direction, see Equation (4.27).
3.4. Calculations of morphisms
Now we calculate the morphisms $HF(\mathbf{L}_{i},\mathbf{L}_{j})$ on the A-side where $\mathbf{L}$ is $L$ or $K$. We start with the differential between two U-shapes $L_{i}$ and $L_{j}$, which involves counting bigons over the one depicted in Figure 5 from $b_{1}$ to $b_{0}=b_{S}$. Note the difference from Figure 4; bigons between U-shapes instead go from degree 1-1 to degree 0 in the base. In a fiber, by [ACLL_tori, §4.2, Equation (4.21)] with $a_{i}=b_{i}=0$, $g=2$,
(3.53) $HF^{d}(\ell_{i},\ell_{j})=\begin{cases}\delta_{d0}\cdot\mathbb{C}^{(i-j)^{2}}&i<j\\ \mathbb{C}^{2\choose d}&i=j\\ \delta_{d2}\cdot\mathbb{C}^{(i-j)^{2}}&i>j\end{cases}$
where $\delta$-function $\delta_{d0}$ is 0 unless $d=0$ in which case it is 1, and similarly for $\delta_{d2}$.
Example 3.24.
$CF^{d}(\ell_{i+1},\ell_{i})=HF^{d}(\ell_{i+1},\ell_{i})$ only has intersection points in degree $d=2$ because the input Lagrangian has larger slope than the output Lagrangian [ACLL_Bfield, Example 4.6].
Then because under monodromy $\Phi(\ell_{i})$ is Hamiltonian isotopic fo $\ell_{i+1}$ [Ca20, Lemma 4.24], $HF^{d}(L_{i},L_{j})$ is defined by the cohomology in degree $d$ of
(3.54) $CF^{d}_{b_{1}}({\ell}_{i+1},{\ell}_{j})[-1]\oplus CF^{d-1}_{b_{0}}({\ell}_{i},{\ell}_{j})\xrightarrow[]{\partial^{d-1}}CF^{d+1}_{b_{1}}({\ell}_{i+1},{\ell}_{j})[-1]\oplus CF^{d}_{b_{0}}(\ell_{i},\ell_{j})$
$\xrightarrow[]{\partial^{d}}CF^{d+2}_{b_{1}}({\ell}_{i+1},{\ell}_{j})[-1]\oplus CF^{d+1}_{b_{0}}(\ell_{i},\ell_{j}).$
$CF^{d}_{b_{1}}({\ell}_{i},{\ell}_{j})[-1]$ means we take the degree $d$ intersection points in the fiber over $b_{1}$ and degree 1-1 in the base. All differentials within a fiber
(3.55) $\partial:CF^{d}_{b}(\ell_{i},\ell_{j})\xrightarrow[]{0}CF^{d+1}_{b}(\ell_{i},\ell_{j})$
are zero. There are no bigons between two linear Lagrangians in the universal cover for $i\neq j$, and when $i=j$ we use [ACLL_tori, Equation (4.47)] for local systems $a_{1}=a_{2}=0$ here. Furthermore, given the location of the Lagrangians, there can be no bigons from $b_{0}$ to $b_{1}$ as they must be orientation preserving. So
(3.56) $\partial:CF_{b_{0}}(\ell_{i},\ell_{j})\xrightarrow[]{0}CF_{b_{1}}(\ell_{i^{\prime}},\ell_{j^{\prime}}).$
As we will see below in the Proof of Corollary 4.3, the possible short exact sequences computing $HF(L_{i},L_{j})$ are therefore the following. Each $\partial$ counts bigons over the base bigon from ${b}_{1}$ to ${b}_{0}$.
(3.57) $0\to CF^{0}_{{b_{1}}}(\ell_{i+1},\ell_{j})[-1]$ $\xrightarrow[]{\partial}CF^{0}_{{b_{0}}}(\ell_{i},\ell_{j})\to HF^{0}({L}_{i},{L}_{j})\to 0,\qquad$ $j>i+1$
$0$ $\to CF^{d+1}_{{b}_{1}}({\ell}_{i+1},{\ell}_{i+1})[-1]=HF^{d}({L}_{i},{L}_{i+1})\to 0,\qquad$ $d=0,1$
$0$ $\to CF^{d}_{{b}_{0}}(\ell_{i},\ell_{i})=HF^{d}({L}_{i},{L}_{i})\to 0,\qquad$ $d=0,1$
$0\to CF^{2}_{{b_{1}}}(\ell_{i+1},\ell_{j})[-1]$ $\xrightarrow[]{\partial}CF^{2}_{{b_{0}}}(\ell_{i},\ell_{j})\to HF^{2}({L}_{i},{L}_{j})\to 0,\qquad$ $j<i$
It remains to calculate the differential in the first line in Equation (3.57) (the differential in the last line is the same). We assume $j>i+1$.
Lemma 3.25 (Morphisms on $\mathcal{A}_{L}$).